OpenStudy (anonymous):

PLEASE HELP! Use the quadratic formula to solve 3x^2 – 4x – 1 = 0. If necessary, round answers to the nearest hundredth.

OpenStudy (anonymous):

i think the answer is 1.55, -0.22

jimthompson5910 (jim_thompson5910):

The equation \(\Large 3x^2 – 4x – 1 = 0 \) is in the form \(\Large ax^2+bx+c = 0\) where a = 3, b = -4 and c = -1 Plug this into the quadratic formula and simplify to solve for x. \[\Large x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\] \[\Large x=\frac{-\left(-4\right)\pm\sqrt{\left(-4\right)^{2}-4\left(3\right)\left(-1\right)}}{2\left(3\right)}\] \[\Large x=\frac{4\pm\sqrt{16-\left(-12\right)}}{6}\] \[\Large x=\frac{4\pm\sqrt{16+12}}{6}\] \[\Large x=\frac{4\pm\sqrt{28}}{6}\] \[\Large x=\frac{4+\sqrt{28}}{6}\ \text{or} \ x=\frac{4-\sqrt{28}}{6}\] \[\Large x=\frac{4+2\sqrt{7}}{6}\ \text{or} \ x=\frac{4-2\sqrt{7}}{6}\] \[\Large x \approx 1.55\ \text{or} \ x \approx -0.22\] So you are absolutely correct. Nice work.

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

sure thin

jimthompson5910 (jim_thompson5910):

thing*

OpenStudy (anonymous):

thank u so much i got an a!!!

jimthompson5910 (jim_thompson5910):

you're very welcome

jimthompson5910 (jim_thompson5910):

I didn't do much this time around since you already got the answers lol

OpenStudy (anonymous):

its ook thank you thank you thank you i have taken this over and over again and i only just finally passed it was 40 questions and i could only miss 4

jimthompson5910 (jim_thompson5910):

oh wow, that's a steep requirement

OpenStudy (anonymous):

haha my school is stupid

jimthompson5910 (jim_thompson5910):

lol sounds like it

jimthompson5910 (jim_thompson5910):

but I'm glad you finally made it

OpenStudy (anonymous):

thank you now i can go to bed talk to you later!

jimthompson5910 (jim_thompson5910):

have a great night's sleep, cya later